# High-dimensional dynamics in low-dimensional networks

### v4

@@ Methods
Acknowledgments
We thank Ashok Litwin-Kumar and David G. Clark for helpful conversations and comments on drafts of the manuscript. This work was supported by the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0223 and the National Science Foundation under a Neuronex award number DBI-1707400.
Appendix
@@ Appendix A Comparison to previous work on dynamics of low-rank networks
In this section we review the relationship between our models and the models from previous theoretical work on low-rank networks, specifically work by Ostojic et al. [2, 6, 4, 7, 8, 9], work by Thibeault et al. [1], and work by Landau and Sompolinsky [5, 3]
@@
where $r$ is the rank and each ${\bf{m}}_{\mu}$ and ${\bf{n}}_{\mu}$ are $N\times 1$ vectors with entries that scale like $\mathcal{O}(1)$. Specifically, they are taken to be random vectors and the variance of their entries scales like $\mathcal{O}(1)$. Often, they are taken to be biased random vectors with a non-zero mean that also scales like $\mathcal{O}(1)$. Because of the $1/N$ factor in Eq. (12), the variance of the entries in $W_{0}$ scale like $\mathcal{O}(1/N^{2})$ and, when the entries are biased, the mean entry in $W_{0}$ scales like $\mathcal{O}(1/N)$. Regardless of whether entries are biased, the singular values of $W_{0}$ scale like $\mathcal{O}(1)$ in the models considered by Ostojic and colleagues [2]. Hence, the singular values of the low-rank part ($W_{0}$) and the random part ($W_{1}$) have the same scale in the work of Ostojic and colleagues, in contrast to our models in which the singular values of $W_{0}$ are considered to be asymptotically larger than the singular values of $W_{1}$. This difference in scaling is acknowledged by Ostojic et al. [2, 6, 4, 7, 8, 9] who refer to their networks as “weakly low-rank.”
The distinction between strongly and weakly low-rank networks is more nuanced when the full-rank, random part, $W_{1}$, of the connectivity matrix is assumed to be weak or absent. For example, recent work by Mastrogiuseppe et al. [10] considered the case in which the full-rank part is absent ($W_{1}=0$) and the low-rank connectivity matrix, $W=W_{0}$, has $\mathcal{O}(1)$ singular values. While their use of $\mathcal{O}(1)$ singular values is similar to weakly low-rank networks from other work [2], the absence of a random part means that their connectivity matrices are truly low-rank. In that work [10], Mastrogiuseppe et al. showed that the emergence of suppression and high-dimensional dynamics depends on the degree of overlap between the left- and right-singular vectors in ways that are not captured by our asymptotic theory.
For the sake of comparison to the models in [10], let us consider a rank-one network structure like the one in Figure 1. Specifically, we take $W_{1}$ to be random with spectral radius $\rho>0$, and take
@@
In Figure 1 and its analysis, we assume that $\rho\sim\mathcal{O}(1)$. In this case, a strong (i.e., asymptotically dominating) low-rank structure requires that $|c|\gg 1$. Therefore, stability in this case requires that $c<0$, which produces strong, negative eigenvalues (as in Figure 1; however, consider also the examples in Figure 2, Figure 3k–m, and Figure 12 in Appendix D which are strongly low-rank without large eigenvalues). We could alternatively take the random, full-rank part to be weak: $\rho\ll 1$ (or it could be absent, $\rho=0$, as in [10]). In this case, we can take the low-rank part to be moderate in magnitude ($|c|=\mathcal{O}(1)$) while maintaining an asymptotically dominant low-rank structure because $|c|\gg\rho$ (Figures 7 and 8 in Appendix D).
Figure 7: Response properties of a weakly low-rank network with $c>0$. Same as Figure 1 except c=0.5c=-0.5 and $\rho=0.01$.
Figure 8: Response properties of a weakly low-rank network with $c<0$. Same as Figure 1 except $c=0.5$ and $\rho=0.01$.
When $c<0$, the network exhibits a weaker form of low-rank suppression along with high-dimensional responses (Figure 7). Similar examples and results are considered in [10]. In conclusion, some of our theoretical results on low-rank suppression and high-dimensional dynamics require that the low-rank part is large in absolute terms ($|c|\gg 1$), not just relative terms ($|c|\gg\rho$).
Since $|c|=\mathcal{O}(1)$, we can now take $c>0$ while keeping the network stable with a dominant low-rank part (in contrast to strongly low rank networks like Figure 1 where stability requires $c<0$). When $c>0$, inputs aligned to $\bf{u}$} are amplified compared to random perturbations (Figure 8), reversing the trend of low-rank suppression. However, the amplification is weak. In response to high-dimensional input, the network dynamics have a dominant principal component direction, but the dominance is weaker than in other examples we have considered (Figure 8d; compare to Figure 2 and Figure 12).
@@
Appendix B Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
@@
We now consider conditions for the emergence of high-dimensional dynamics (as in Figure 1d). We assume that ${\bf{x}}(t)$ is a stationary, ergodic stochastic process, as in Figure 1d–f. Define the cross-spectral matrix, $\widetilde{C}^{{\bf{x}}}(f)$, of ${\bf{x}}(t)$ at frequency $f$ as the Fourier transform of the matrix of cross-covariance matrix,
@@
where $C_{jk}^{{\bf{x}}}(\tau)=\text{cov}({\bf{x}}_{j}(t),{\bf{x}}_{k}(t+\tau))$ is the stationary cross-covariance. When each ${\bf{x}}_{j}(t)$ is i.i.d., then $\widetilde{C}^{{\bf{x}}}(f)=I\widetilde{a}_{x}(f)$ is a multiple of the identity matrix where $\widetilde{a}_{x}(f)$ is the power spectral density of each ${\bf{x}}_{j}(t)$. The assumption of i.i.d. elements is true for all of the examples we consider, but we will not apply this simplification until the end of our calculation.
The the cross-spectral density of the network response, ${\bf{z}}(t)$, is defined analogously and it can be derived under the dynamics Eq. (6) to get [31, 53, 41, 42, 54]
$\widetilde{C}^{{\bf{z}}}(f)$ $=[\widetilde{r}(f)I-W]^{-1}\widetilde{C}^{{\bf{x}}}(f)[\widetilde{r}(f)I-W]^{-*}$ (13)
@@
$\overline{C}^{{\bf{z}}}_{jk}=\text{cov}({\bf{z}}_{j}(t),{\bf{z}}_{k}(t))=\int_{-\infty}^{\infty}\widetilde{C}^{{\bf{z}}}_{jk}(f)df.$
@@
$\overline{C}^{{\bf{z}}}=\int_{-\infty}^{\infty}[\widetilde{r}(f)I-W]^{-1}\widetilde{C}^{{\bf{x}}}(f)[\widetilde{r}(f)I-W]^{-*}df$ (14)
If the timescale of fluctuations in ${\bf{x}}(t)$ are much slower than the timescale, $\tau$, of network dynamics then $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>\epsilon/\tau$ where $\epsilon\ll 1$. Indeed, we can take this to be the definition of the statement that the fluctuations in ${\bf{x}}(t)$ are much slower than $\tau$. Since $\widetilde{r}(f)\approx 1$ whenever $f\tau<\epsilon$, we therefore have that $\widetilde{r}(f)\approx 1$ whenever $\widetilde{C}^{{\bf{x}}}(f)$ is not close to zero. As a result, the only parts of the integrand that contribute to the integral in Eq. (14) are the low frequency components, $f\approx 0$. In this case, we can replace $\widetilde{r}(f)$ with $1$ to obtain
@@
Note that Eq. (15) can also be derived directly from the quasi-steady state approximation ${\bf{z}}(t)\approx[I-W]^{-1}{\bf{x}}(t)$. Our assumption above that $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>\epsilon/\tau$ is essentially equivalent to the assumption that the quasi-steady state approximation is accurate, but our approach here is more mathematically precise.
@@
In more general classes of models in which $\overline{C}^{{\bf{x}}}$is not a multiple of the identity matrix, the dimensionality of ${\bf{z}}(t)$ might be reduced when $\overline{C}^{{\bf{x}}}$ is effectively low-rank. However, this effect would be caused by a low-dimensional perturbation ($\overline{C}^{{\bf{x}}}$ low-dimensional) (see Figure 11), and would not be related to the intrinsic dynamics of interactions within the network.
@@
More specifically, the variance explained by each principal component of ${\bf{z}}(t)$ is given by the square of the singular values of $A$,
@@
To demonstrate these analytical results, we repeated the simulation from Figure 1d and added the predictions from Eq. (16) as red dots (Figure 9a). Surprisingly, the theory did not closely match the simulations. We suspected that this was due to finite sampling: The simulation was performed over the time interval $t\in[0,T]$ where $T=5\times 10^{3}$ (for comparison, $\tau=1$ and the correlation timescale of ${\bf{x}}(t)$ was $\tau_{x}=5$). We suspected that Eq. (15) would be accurate when $\overline{C}^{\bf{x}}$ is replace by the empirical covariance matrix of ${\bf{x}}(t)$. Under this substitution, Eq. (16) would be replaced by
$\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\frac{1}{\sigma^{2}_{U,N-k}}$
@@
$U=[I-W]\sqrt{{\overline{C}}^{\bf{x}}}^{-1}$
and the empirical covariance matrix is used for ${\overline{C}}^{\bf{x}}$. Or, equivalently and more simply,
@@
$\hat{R}=[I-W]^{-1}\sqrt{{\overline{C}}^{\bf{x}}}$
is the sampled value of $R$ and the empirical value of ${\overline{C}}^{\bf{x}}$ is again used. Using Eq. (17) gives a much more accurate prediction (Figure 9a, green dots). This confirms that the error in the red dots from Figure 9a is due largely to under-sampling of ${\bf{x}}(t)$. We next increased the simulation time ten-fold to $T=5\times 10^{4}$. In this case, the original Eq. (16) was more accurate (Figure 9b, red dots), further confirming that the errors in Figure 9a are due largely to sampling error.
As concluded previously, ${\bf{z}}(t)$ exhibits low-rank suppression (as demonstrated by the last blue dot in Figure 1d) whenever $Q$ has a small number of asymptotically large singular values (equivalently, whenever $A$ has a small number of asymptotically small singular values). Now, we may conclude that ${\bf{z}}(t)$ has low-dimensional dynamics whenever $Q$ has some asymptotically small singular values (equivalently, whenever $A$ has a small number of asymptotically large singular values). Conversely, high-dimensional dynamics can only occur whenever $Q$ does not have any asymptotically small singular values (equivalently, whenever $A$ does not have any asymptotically large singular values).
Combining these conclusions with those in Claim 1, we can summarize as follows
@@
Previously, in proving Claim 1, we effectively showed that Q=IWQ=I-W has large singular values (and therefore low-rank suppression occurs) whenever $W$ is effectively low-rank. Conditions under which Q=IWQ=I-W lacks small singular values (and therefore high-dimensional dynamics occur) are not so simple. Specifically, the lack or presence of small singular values depends on the recurrent alignment matrix,
@@
The dynamics of ${\bf{z}}(t)$ are low dimensional whenever ${\bf{z}}(t)$ has a small number of dominant principal components as in Figure 2 and Figure 3g (conversely, ${\bf{z}}(t)$ is high-dimensional whenever there is no such dominant principal component, as in Figure 1d). From the discussion above, we know that ${\bf{z}}(t)$ is low dimensional whenever $A=[I-W]^{-1}$ has at least one asymptotically large singular value or, equivalently, whenever Q=IWQ=I-W has an asymptotically small singular value. Therefore, our original claim is equivalent to the following:

### v5

@@ IX Methods
Acknowledgements.
We thank Ashok Litwin-Kumar and David G. Clark for helpful conversations and comments on drafts of the manuscript. This work was supported by the Air Force Office of Scientific Research (AFOSR) awards numbers FA9550-21-1-0223 and FA9550-26-1-0004, and the National Science Foundation under a Neuronex award number DBI-1707400.
@@ Appendix A Comparison to previous work on dynamics of low-rank networks
Here, we focus on the response of networks to high-dimensional external input. Some previous work on low-rank networks and low-dimensional dynamics focuses on the dimensionality of intrinsic dynamics generated by the network in the absence of external input or in the presence of low-dimensional external input. For example chaotic and other nonlinear dynamics in untrained and trained networks can be low-dimensional (see [22, 21, 2, 23, 1] and Figures 10 and 11 in Appendix F). In contrast, we are here primarily interested in the network’s response to high-dimensional external input. In our model’s, the network’s activity is driven primarily by the external input (which is filtered by the network’s internal structure) and not by intrinsic dynamics. When the network is strongly low-rank, we showed that low-dimensional dynamics can only arise when the recurrent alignment matrix has small singular values (see Claim 1).
In this section we review the relationship between our models and the models from previous theoretical work on low-rank networks, especially work by Ostojic et al. [3, 7, 5, 8, 9, 10], work by Thibeault et al. [1], and work by Landau and Sompolinsky [6, 4]
@@
where $r$ is the rank and each ${\bf{m}}_{\mu}$ and ${\bf{n}}_{\mu}$ are $N\times 1$ vectors with entries that scale like $\mathcal{O}(1)$. Specifically, they are taken to be random vectors and the variance of their entries scales like $\mathcal{O}(1)$. Often, they are taken to be biased random vectors with a non-zero mean that also scales like $\mathcal{O}(1)$. Because of the $1/N$ factor in Eq. (10), the variance of the entries in $W_{0}$ scale like $\mathcal{O}(1/N^{2})$ and, when the entries are biased, the mean entry in $W_{0}$ scales like $\mathcal{O}(1/N)$. Regardless of whether entries are biased, the singular values of $W_{0}$ scale like $\mathcal{O}(1)$ in the models considered by Ostojic and colleagues [3]. Hence, the singular values of the low-rank part ($W_{0}$) and the random part ($W_{1}$) have the same scale whenever the spectral radius of $W_{1}$ is $\mathcal{O}(1)$, in contrast to our models in which the singular values of $W_{0}$ are considered to be asymptotically larger than the singular values of $W_{1}$. This difference in scaling is acknowledged by Ostojic et al. [3, 7, 5, 8, 9, 10] who refer to their networks as “weakly low-rank.”
To summarize, when $W_{1}$ is a random matrix with $\mathcal{O}(1)$ spectral radius and when the entries of $W_{0}$ are biased, the primary difference between many strongly and weakly low-rank networks in practice is that the entries of $W_{0}$ scale like $\mathcal{O}(1/N)$ in weakly low rank networks while they are larger (typically $\mathcal{O}(1/\sqrt{N})$) in many strongly low-rank networks, for example in work on balanced networks [6, 4]. The transition between these two regimes can be visualized by gradually scaling the magnitude of the low-rank part from weak to strong. Figure 22 in Appendix F shows that low-rank suppression arises gradually in this case.
The distinction between strongly and weakly low-rank networks is more nuanced when the full-rank, random part, $W_{1}$, of the connectivity matrix is assumed to be weak or absent. For example, recent work by Mastrogiuseppe et al. [11] considered the case in which the full-rank part is absent ($W_{1}=0$) and the low-rank connectivity matrix, $W=W_{0}$, has $\mathcal{O}(1)$ singular values. While their use of $\mathcal{O}(1)$ singular values is similar to weakly low-rank networks from other work [3], the absence of a random part means that their connectivity matrices are truly low-rank (not just approximately low-rank). In that work [11], Mastrogiuseppe et al. showed that the emergence of suppression and high-dimensional dynamics depends on the degree of overlap between the left- and right-singular vectors in ways that are not captured by our asymptotic theory.
For the sake of comparison to the models in [11], we consider a rank-one network structure like the one in Figure 1. Specifically, we take $W_{1}$ to be random with spectral radius $\rho>0$, and take
@@
In Figure 1 and its analysis, we assume that $\rho\sim\mathcal{O}(1)$. In this case, a strong (i.e., asymptotically dominating) low-rank structure requires that $|c|\gg 1$. Therefore, stability in this case requires that $c<0$, which produces strong, negative eigenvalues (as in Figure 1; however, consider also the examples in Figure 2, Figure 3k–m, and Figure 9 which are strongly low-rank without large eigenvalues). We could alternatively take the random, full-rank part to be weak: $\rho\ll 1$ (or it could be absent, $\rho=0$, as in [11]). In this case, we can take the low-rank part to be moderate in magnitude ($|c|=\mathcal{O}(1)$) while maintaining an asymptotically dominant low-rank structure because $|c|\gg\rho$ when $|c|=\mathcal{O}(1)$ and $\rho\ll 1$. Repeating the simulation from Figure 1 demonstrates a weakened form of low-rank suppression with high-dimensional dynamics when $c<0$ and an amplification of aligned inputs with weakly low-dimensional dynamics when $c>0$ (Figures 20 and 21 in Appendix F). We also repeated all of the simulations from Figure 3 in networks with $\rho\ll 1$ and $\sigma_{1},\sigma_{2}\sim\mathcal{O}(1)$, so that the network is still low-rank in the sense that the first two singular values dominate, but not strongly low-rank because they are $\mathcal{O}(1)$ (see Figure 23a). Under these conditions, the dynamics were high-dimensional regardless of the structure of $W_{0}$ (Figure 23b–f). To understand this result, note that low-dimensional dynamics in our models (Figures 2 and 3i,l) arise from strong amplification of a small number of directions by $W$. When the low-rank part of $W$ is weaker, these low-dimensional dynamics will be weaker or absent. In other words, when the dominant singular values of $W_{0}$ are $\mathcal{O}(1)$ and smaller, the identity in $I-W]^{-1}$ from the quasi-stead state analysis more strongly dominates the $W$ (or, at least, the $W$ does not dominate the $I$). More directly, the ${\bf{x}}(t)$ in Eq. (1) dominates the $W{\bf{z}}$. Therefore, high-dimensional inputs are simply transferred to high-dimensional responses when $W$ is weaker.
When $c<0$, the network exhibits a weaker form of low-rank suppression along with high-dimensional responses (Figure 20 in Appendix F). Similar examples and results are considered in [11]. In conclusion, some of our theoretical results on low-rank suppression and high-dimensional dynamics require that the low-rank part is large in absolute terms ($|c|\gg 1$), not just relative terms ($|c|\gg\rho$).
When $|c|=\mathcal{O}(1)$, we can take $c>0$ while keeping the network stable with a dominant low-rank part (in contrast to strongly low rank networks like Figure 1 where stability requires $c<0$). When $c>0$, inputs aligned to $\bf{u}$} are amplified compared to random perturbations (Figure 21 in Appendix F), reversing the trend of low-rank suppression. However, the amplification is weak. In response to high-dimensional input, the network dynamics have a dominant principal component direction, but the dominance is weaker than in other examples we have considered (Figure 21d; compare to Figure 2 and Figure 9) because stability imposes a bound on $c$ (specifically, $c<1$) and therefore on $\|W_{0}\|$.
@@
Appendix B Details for the linearization of dynamics around a stable equilibrium
Here, we make precise our assumption that dynamics represent a small perturbation around a stable equilibrium, and give a more precise derivation of the linearized dynamics.
We begin by considering nonlinear dynamics of the form
$\tau\frac{d{\bf{z}}}{dt}={\bf{F}}({\bf{z}},\boldsymbol{\eta})$
where $\boldsymbol{\eta}(t)$ is the external input before any change of coordinates is applied. We assume that there exists a constant-in-time, $\boldsymbol{\eta}(t)=\boldsymbol{\eta}_{0}$, that produces a stable steady-state solution, ${\bf{z}}_{0}$. In other words, there exist $\boldsymbol{\eta}_{0}$ and ${\bf{z}}_{0}$ such that
${\bf{F}}({\bf{z}}_{0},\boldsymbol{\eta}_{0})={\bf{z}}_{0}.$
$J=-I+\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\boldsymbol{\eta}_{0})$
has eigenvalues with strictly negative real part. Here, $I$ is the identity matrix and $\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\boldsymbol{\eta}_{0})$ is the Jacobian matrix of $\bf{F}$} with respect to $\bf{z}$} evaluated at the fixed point. We then consider a small perturbation around this fixed point driven by a perturbation to the forcing term,
$\boldsymbol{\eta}_{p}(t)=\boldsymbol{\eta}_{0}+\epsilon\boldsymbol{\eta}(t).$
The response, ${\bf{z}}_{p}(t)$, of the network to the perturbed forcing term, $\boldsymbol{\eta}_{p}(t)$, can be written to linear order in $\epsilon$ as
${\bf{z}}_{p}(t)={\bf{z}}_{0}+\epsilon{\bf{z}}(t)+\mathcal{O}(\epsilon^{2}).$
The perturbation, ${\bf{z}}(t)$, obeys the linearized equation
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+{\bf{x}}(t)$ (11)
where
$W=\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\boldsymbol{\eta}_{0})$
is interpreted as the effective connectivity matrix and
${\bf{x}}(t)=\partial_{\boldsymbol{\eta}}{\bf{F}}({\bf{z}}_{0},\boldsymbol{\eta}_{0})\boldsymbol{\eta}(t)$
is interpreted as the effective input. Our assumption that the equilibrium is stable implies that all eigenvalues of $W$ have real part less than $1$. We additionally assume that the real part of all eigenvalues of the Jacobian are bounded away from 0, meaning that the real part of all eigenvalues of $W$ are bounded below $1$. In other words, there is a $d>0$ with $d\sim\mathcal{O}(1)$ such that all eigenvalues of $W$ satisfy $\textrm{Re}(\lambda_{W})<1-d$. In particular, this implies that $\rho<1-d$ where $\rho$ is the spectral radius of the random part, $W_{1}$, of $W$.
Note that Eq. (11) is equivalent to the linear dynamics in Eq. (1). Hence, the linear network dynamics in Eq. (1) are a sort of normal form for the linearized dynamics described here. Therefore, we are justified in focusing on the linear dynamics of Eq. (1) in the analysis in the Results.
In many models, the perturbation is purely additive, ${\bf{F}}({\bf{z}},\boldsymbol{\eta})={\bf{H}}({\bf{z}})+\boldsymbol{\eta}$, so that the effective input is equal to the raw input, ${\bf{x}}=\boldsymbol{\eta}$. In other models, the raw input can have a different structure and dimensionality to the effective input. For illustrative purposes, we consider a contrived example of such a model in Figure 8 in Appendix F.
Some models are parameterized in terms of a connectivity matrix that is not equal to the Jacobian matrix. For example, the system might be written as $d{\bf{z}}/dt={\bf{G}}(W{\bf{z}},{\bf{x}})$ where $W$ is meant to quantify a connectivity matrix, but $\partial_{\bf{z}}{\bf{G}}(W{\bf{z}}_{0},{\bf{x}}_{0})\neq W$. In this case, our interpretation of the effective connectivity matrix is not consistent with the original parameterization, but they will often share structural properties such as their approximate dimensionality. We consider one such example from epidemiology at the end of the Results. In that example, we found that the use of effective connectivity in place of the explicit connectivity matrix is important for our results to hold (see Figure 19 in Appendix F)
Appendix C Precise definitions of strongly low-rank networks and low-dimensional dynamics.
As mentioned above, we focus our attention on strongly low-rank networks. Intuitively, this means that the connectivity matrix has a small number of large singular values and the rest are not large. To make this definition mathematically precise, we need to consider an asymptotic limit of large $N$, where $N$ is the network size. Let $\sigma_{k}$ be the $k$th largest singular value of $W$. In strongly low-rank networks, there exists an $r\sim\mathcal{O}(1)$ such that $\lim_{N\to\infty}\sigma_{k}=\infty$ for $k=1,\ldots,r$. In addition, $\sigma_{k}\leq\mathcal{O}(1)$ for $k>r$.
We additionally need to specify what we mean by “low-dimensional” and “high-dimensional” dynamics. Let $\lambda_{k}$ be the $k$th largest eigenvalue of the covariance matrix of ${\bf{z}}(t)$, equivalently the $k$th largest singular value since covariance matrices are symmetric and positive definite. Note that $\lambda_{k}$ is a non-negative real number representing the amount of variance explained by the $k$th principal component of ${\bf{z}}(t)$. Intuitively, ${\bf{z}}(t)$ is low-dimensional whenever a large portion of the variance is explained by a small number of singular values. To make this precise, we say that the dynamics of ${\bf{z}}(t)$ are low-dimensional whenever there is an $r\sim\mathcal{O}(1)$ such that $\lim_{N\to\infty}\lambda_{j}/\lambda_{k}=0$ whenever $j>r$ and $k\leq r$.
@@ Appendix B Details for the linearization of dynamics around a stable equilibrium
Appendix D Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
@@
We now consider conditions for the emergence of high-dimensional dynamics. We assume that ${\bf{x}}(t)$ is a stationary, ergodic stochastic process. Define the cross-spectral matrix, $\widetilde{C}^{{\bf{x}}}(f)$, of ${\bf{x}}(t)$ at frequency $f$ as the Fourier transform of the matrix of cross-covariance matrix,
@@
where $C_{jk}^{{\bf{x}}}(\tau)=\text{cov}({\bf{x}}_{j}(t),{\bf{x}}_{k}(t+\tau))$ is the stationary cross-covariance. When each ${\bf{x}}_{j}(t)$ is i.i.d., then $\widetilde{C}^{{\bf{x}}}(f)=Ic_{x}(f)$ is a multiple of the identity matrix where $c_{x}(f)$ is the power spectral density of each ${\bf{x}}_{j}(t)$. The assumption of i.i.d. elements is true for all of the examples we consider, but we will not apply this simplification until later in our calculation.
The the cross-spectral density of the network response, ${\bf{z}}(t)$, is defined analogously and it can be derived under the dynamics in Eq. (11) to get [42, 59, 40, 41, 60]
$\widetilde{C}^{{\bf{z}}}(f)$ $=[r(f)I-W]^{-1}\widetilde{C}^{{\bf{x}}}(f)[r(f)I-W]^{-*}$ (12)
@@
$\overline{C}^{{\bf{z}}}_{jk}=\text{cov}({\bf{z}}_{j}(t),{\bf{z}}_{k}(t))=\int_{-\infty}^{\infty}\widetilde{C}^{{\bf{z}}}_{jk}(f)df.$ (13)
@@
$\overline{C}^{{\bf{z}}}=\int_{-\infty}^{\infty}[r(f)I-W]^{-1}\widetilde{C}^{{\bf{x}}}(f)[r(f)I-W]^{-*}df$ (14)
If the timescale of fluctuations in ${\bf{x}}(t)$ are much slower than the timescale, $\tau$, of network dynamics then $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ at frequencies much higher than $1/(2\pi\tau)$ (see Figure 13 in Appendix F) . Indeed, we can take this to be the definition of the statement that the fluctuations in ${\bf{x}}(t)$ are much slower than $\tau$. Since $r(f)\approx 1$ whenever $f\ll 1/(2\pi\tau)$, we therefore have that $r(f)\approx 1$ whenever $\widetilde{C}^{{\bf{x}}}(f)$ is not close to zero (see Figure 13 in Appendix F). As a result, the only parts of the integrand that contribute to the integral in Eq. (14) are the low frequency components, $f\approx 0$. In this case, we can replace $r(f)$ with $1$ to obtain
@@
Note that Eq. (15) can also be derived directly from the quasi-steady state approximation ${\bf{z}}(t)\approx[I-W]^{-1}{\bf{x}}(t)$. Our assumption above that $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>1/\tau$ is essentially equivalent to the assumption that the quasi-steady state approximation is accurate. Either approach can be used to derive Eq. (15). In Appendix E below, we provide more details about the frequency dependence of variability and how it relates to the speed at which ${\bf{x}}(t)$ varies.
@@
In more general classes of models in which $\overline{C}^{{\bf{x}}}$is not a multiple of the identity matrix, the dimensionality of ${\bf{z}}(t)$ might be reduced when $\overline{C}^{{\bf{x}}}$ is effectively low-rank (see Figure 8 in Appendix F). However, in this situation, low-dimensional dynamics are inherited from the input, not generated by interactions within the network. By restricting to high-dimensional ${\bf{x}}(t)$, we are focused on the question of whether low-dimensional dynamics arise from network interactions.
@@
More specifically, the variance explained by each principal component of ${\bf{z}}(t)$ is proportional to the square of the singular values of $A$,
@@
To demonstrate these analytical results, we repeated the simulation from Figure 1d and added the predictions from Eq. (16) as red dots (Figure 24a in Appendix F). Surprisingly, the theory did not closely match the simulations. We suspected that this was due to finite sampling: The simulation was performed over the time interval $t\in[0,T]$ where $T=5\times 10^{3}$ (for comparison, $\tau=1$ and the correlation timescale of ${\bf{x}}(t)$ was $\tau_{x}=10$). We suspected that Eq. (15) would be accurate when $\overline{C}^{\bf{x}}$ is replace by the empirical covariance matrix of ${\bf{x}}(t)$. Under this substitution, Eq. (16) would be replaced by
$\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\frac{1}{\sigma^{2}_{\hat{B},N-k}}$
@@
$\hat{B}=[I-W]\sqrt{{\hat{C}}^{\bf{x}}}^{-1}$
and ${\hat{C}}^{\bf{x}}$ is the the empirical covariance matrix. Or, equivalently and more simply,
@@
$\hat{R}=[I-W]^{-1}\sqrt{{\hat{C}}^{\bf{x}}}$
is the sampled value of $R$. Using Eq. (17) gives a more accurate prediction (Figure 24a in Appendix F, green dots). This confirms that the error in the red dots from Figure 24a is due largely to under-sampling of ${\bf{x}}(t)$. We next increased the simulation time ten-fold to $T=5\times 10^{4}$. In this case, the original Eq. (16) was more accurate (Figure 24b in Appendix F, red dots), further confirming that the errors in Figure 24a are due largely to sampling error.
As concluded previously, ${\bf{z}}(t)$ exhibits low-rank suppression (as demonstrated by the last blue dot in Figure 1d) whenever $Q$ has a small number of asymptotically large singular values (equivalently, whenever $A$ has a small number of asymptotically small singular values). Now, we may conclude that ${\bf{z}}(t)$ has low-dimensional dynamics whenever $Q$ has some asymptotically small singular values (equivalently, whenever $A$ has a small number of asymptotically large singular values). Conversely, high-dimensional dynamics can only occur whenever $Q$ does not have any asymptotically small singular values (equivalently, whenever $A$ does not have any asymptotically large singular values). We can summarize as follows
@@
Previously, in Section IV.5, we effectively showed that Q=IWQ=I-W has large singular values (and therefore low-rank suppression occurs) whenever $W$ is strongly low-rank. Conditions under which Q=IWQ=I-W lacks small singular values (and therefore high-dimensional dynamics occur) are not so simple. Specifically, the lack or presence of small singular values depends on the recurrent alignment matrix,
@@
The dynamics of ${\bf{z}}(t)$ are low dimensional whenever ${\bf{z}}(t)$ has a small number of dominant principal components as in Figure 2d (conversely, ${\bf{z}}(t)$ is high-dimensional whenever there is no such dominant principal component, as in Figure 1d). From the discussion above, we know that ${\bf{z}}(t)$ is low dimensional whenever $A=[I-W]^{-1}$ has at least one asymptotically large singular value or, equivalently, whenever Q=IWQ=I-W has an asymptotically small singular value. Therefore, our original claim is equivalent to the following: